IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

Serge Chastel

Heather Flewelling

  • at panstarrs booth at open house
  • addstar restart/stop during shutdowns
  • cleanup of space on ipp005 - accomplished by moving LAP.slice12.% from ipp005 to ipp007, adjusting the database, and cleaning up on ipp005
  • skyprobe datastore investigations continue, with emails to Steph/Haydn
  • investigations of new SAS13 on psps - updated various wikis with results
  • LAP addstar status:
    • queued skycal for all slices we are currently working on - this picks up the few that have been fixed since I queued them initially. These are all addstarred and ready
    • merged slice 12/14/16 into big dvodb
    • addstar is continuing: slice 10/12/14/16/18/22/00 are done, slice 02/20 are still munching
    • slice18 running delstar on this - still in progress (slow...)
    • accidentally did something stupid the other day and had the same label in 2 addstars. I cleaned up this mess

Mark Huber

  • misc liaison, TSS and ops monitoring
  • deepstack tests
  • diffim dual conv tests for MOPS
  • md.pv1 running

Bill Sweeney

Chris Waters